OpenStudy (anonymous):

Find the acute angle between the lines: y=2x+4 and y=-3x+6

OpenStudy (anonymous):

the angle, theta, between lines with known slopes, m1 and m2, is given by the relation: \(\huge tan\theta=\frac{m_2-m_1}{1+m_1m_2} \)

OpenStudy (anonymous):

So I would do the inverse tan of that?

OpenStudy (anonymous):

do you mind to explain how the formula is arrived at?

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

ok... do you know about inclination??

OpenStudy (anonymous):

rise over run?

OpenStudy (anonymous):

no... inclination is an ANGLE a line makes with the positive x axis....

OpenStudy (anonymous):

yes, so that is tan-1(rise/run)

OpenStudy (anonymous):

if in first quadrant

OpenStudy (anonymous):

it doesn't matter if it's in the first quadrant or not.... the inclination is just the angle it makes with the positive x axis at the point where the line intesects the x axis:

OpenStudy (anonymous):

on my calc i was getting negative answers that i had to change around...

OpenStudy (anonymous):

so to get the angle between two lines, you're basically subtracting the inclination of the two lines....

OpenStudy (anonymous):

oh yeah.... sorry, my bad.... the actual formula is: \(\large tan\theta=|\frac{m_2-m_1}{1+m_2m_1}| \)

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

 the angle between the two lines is alpha and you can just use angle sum formula for tangent.